OpenStudy (anonymous):

Simplify the expression The root of negative nine over the quantity of three minus two i plus the quantity of one plus five i. .

OpenStudy (michele_laino):

please can you write your expression using the editor?

OpenStudy (anonymous):

\[(\sqrt{-9})\div(3-2i) + (1+5i)\]

OpenStudy (michele_laino):

here we can rewrite your expression as follows: \[\Large \frac{{\sqrt { - 9} }}{{3 - 2i}} + 1 + 5i = \frac{{3i\left( {3 + 2i} \right)}}{{9 + 4}} + 1 + 5i\]

OpenStudy (michele_laino):

@jordynhazan1

OpenStudy (anonymous):

so that simplifies to \[\frac{ 9i+6i }{ 13 } +1 +5i\]

OpenStudy (anonymous):

right?

OpenStudy (michele_laino):

no, the simplified expression is: \[\Large \begin{gathered} \frac{{\sqrt { - 9} }}{{3 - 2i}} + 1 + 5i = \frac{{3i\left( {3 + 2i} \right)}}{{9 + 4}} + 1 + 5i = \hfill \\ \hfill \\ = \frac{{9i - 6}}{{13}} + 1 + 5i \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

am I right?

OpenStudy (michele_laino):

please remember that: \[\Large {i^2} = - 1\]

OpenStudy (michele_laino):

furthermore, I have used this identity: \[\Large \frac{1}{z} = \frac{{{z^*}}}{{{{\left| z \right|}^2}}}\] where z* is the complex conjugate of z

OpenStudy (anonymous):

i still have to solve for a number but don't know how to

OpenStudy (anonymous):

how do i add the 1 and the 5i to the equation

OpenStudy (michele_laino):

we have to compute the least common multiple, which is 13, so we can write: \[\Large \begin{gathered} \frac{{\sqrt { - 9} }}{{3 - 2i}} + 1 + 5i = \frac{{3i\left( {3 + 2i} \right)}}{{9 + 4}} + 1 + 5i = \hfill \\ \hfill \\ = \frac{{9i - 6}}{{13}} + 1 + 5i = \frac{{9i - 6 + 13\left( {1 + 5i} \right)}}{{13}} = ... \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

please simplify

OpenStudy (michele_laino):

hint: \[\Large \begin{gathered} \frac{{\sqrt { - 9} }}{{3 - 2i}} + 1 + 5i = \frac{{3i\left( {3 + 2i} \right)}}{{9 + 4}} + 1 + 5i = \hfill \\ \hfill \\ = \frac{{9i - 6}}{{13}} + 1 + 5i = \frac{{9i - 6 + 13\left( {1 + 5i} \right)}}{{13}} = \hfill \\ \hfill \\ = \frac{{9i - 6 + 13 + 65i}}{{13}} = ... \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

then it becomes \[\frac{ 74i + 7}{ 13 }\]

OpenStudy (michele_laino):

ok! that's right!

OpenStudy (anonymous):

but that isn't near the answers given

OpenStudy (michele_laino):

Please wait I'm checking my computation

OpenStudy (michele_laino):

sorry your starting expression, at initial post, is different

OpenStudy (michele_laino):

no problem we restart to compute the right expression

OpenStudy (michele_laino):

here is the first step: \[\Large \frac{{\sqrt { - 9} }}{{3 - 2i + 1 + 5i}} = \frac{{3i}}{{4 + 3i}}\]

OpenStudy (michele_laino):

now we can write this: \[\Large \frac{1}{{4 + 3i}} = \frac{{4 - 3i}}{{16 + 9}}\] since I have used this identity: \[\Large \frac{1}{z} = \frac{{{z^*}}}{{{{\left| z \right|}^2}}}\]

OpenStudy (anonymous):

the bottom simplifies to 25

OpenStudy (michele_laino):

ok! substituting into the above expression, we get: \[\Large \begin{gathered} \frac{{\sqrt { - 9} }}{{3 - 2i + 1 + 5i}} = \frac{{3i}}{{4 + 3i}} = \hfill \\ \hfill \\ = \frac{{3i\left( {4 - 3i} \right)}}{{16 + 9}} \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

the answer is C!!

OpenStudy (michele_laino):

or: \[\Large \begin{gathered} \frac{{\sqrt { - 9} }}{{3 - 2i + 1 + 5i}} = \frac{{3i}}{{4 + 3i}} = \hfill \\ \hfill \\ = \frac{{3i\left( {4 - 3i} \right)}}{{16 + 9}} = \frac{{3i\left( {4 - 3i} \right)}}{{25}} = ... \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

\[\frac{ 9 + 12i }{ }\]

OpenStudy (michele_laino):

that's right! It is option C

OpenStudy (anonymous):

over 25

OpenStudy (anonymous):

thank you so much

OpenStudy (michele_laino):

:)